λλA particle falls from rest in a medium in which the resistance is λv^2 per unit mass, v being velocity of the particle at time t. Prove that the distance fallen in time to is 1/λ[Incos(t(√gλ))],where g is the acceleration due to gravity
1
Expert's answer
2020-10-19T18:21:57-0400

Let "m" be the mass of the particle. The weight of the particle is "mg" and the drag force is "m\lambda v^2" .

The equation of motion of the particle is

"m\frac{dv}{dt}=mg-m\lambda v^2\\
\Rightarrow \frac{dv}{dt}=g-\lambda v^2\\
\Rightarrow \frac{dv}{dt}=g(1-\frac{\lambda }{g}v^2)\\
\Rightarrow \frac{dv}{1-\frac{\lambda }{g}v^2}=gdt"

Integrating both sides,

"\int \frac{dv}{1-\frac{\lambda }{g}v^2}=g\int dt+C_1\\
\int \frac{dv}{1-\frac{\lambda }{g}v^2}=gt+C_1\ \quad ......(1)"

where "C_1" is the constant of integration.

Let "z=\sqrt{\frac{\lambda}{g}}\ v\Rightarrow dz=\sqrt{\frac{\lambda}{g}}\ dv"

From (1), "\sqrt{\frac{g}{\lambda}}\int \frac{dz}{1-z^2}=gt+C_1"

"\therefore \sqrt{\frac{g}{\lambda}}\ \tanh^{-1}(z)=gt+C_1" since "\int \frac{dz}{1-z^2}=\tanh^{-1}(z)"

Substituting "z=\sqrt{\frac{\lambda}{g}}\ v" ,

"\sqrt{\frac{g}{\lambda}}\ \tanh^{-1}\left(\sqrt{\frac{\lambda}{g}}\ v\right)=gt+C_1\ \quad .......(2)"

At "t=0, v=0"

"\Rightarrow \sqrt{\frac{g}{\lambda}}\ \tanh^{-1}(0)=0+C_1\\
\Rightarrow C_1=0" since "\tanh^{-1}(0)=0"

From (2),

"\sqrt{\frac{g}{\lambda}}\ \tanh^{-1}\left(\sqrt{\frac{\lambda}{g}}\ v\right)=gt\\
\Rightarrow \tanh^{-1}\left(\sqrt{\frac{\lambda}{g}}\ v\right)=gt\sqrt{\frac{\lambda}{g}}\\
\Rightarrow \tanh^{-1}\left(\sqrt{\frac{\lambda}{g}}\ v\right)=\sqrt{g\lambda }\ t\\
\Rightarrow \sqrt{\frac{\lambda}{g}}\ v=\tanh(\sqrt{g\lambda }\ t)\\
\Rightarrow v=\sqrt{\frac{g}{\lambda}}\tanh(\sqrt{g\lambda }\ t)\\
\Rightarrow \frac{dx}{dt}=\sqrt{\frac{g}{\lambda}}\tanh(\sqrt{g\lambda }\ t)\\
\Rightarrow dx=\sqrt{\frac{g}{\lambda}}\tanh(\sqrt{g\lambda }\ t)dt\\"

Integrating,

"x=\sqrt{\frac{g}{\lambda}}\int \tanh(\sqrt{g\lambda }\ t)dt+C_2\ \quad .......(3)"

where "C_2" is the integration constant.

Let "y=\sqrt{g\lambda }\ t\Rightarrow dy=\sqrt{g\lambda }\ dt"

From (3),

"x=\sqrt{\frac{g}{\lambda}}\ \frac{1}{\sqrt{g\lambda }}\int \tanh(y)dy+C_2\\
\Rightarrow x=\frac{1}{\lambda} \int \tanh(y)dy+C_2\\
\Rightarrow x=\frac{1}{\lambda} \ln|\cosh(y)|+C_2" since "\int \tanh(y)dy=\ln |\cosh(y)|"

Substituting "y=\sqrt{g\lambda }\ t" ,

"x=\frac{1}{\lambda} \ln|\cosh(\sqrt{g\lambda }\ t)|+C_2\ \quad .......(4)"

At "t=0, x=0"

"\Rightarrow 0=\frac{1}{\lambda} \ln|\cosh(0)|+C_2"

Since "\cosh(0)=1 \ and\ \ln(1)=0"

"\Rightarrow C_2=0"

From (4),

"x=\frac{1}{\lambda} \ln|\cosh(\sqrt{g\lambda }\ t)|"

Thus, the distance fallen in time "t" is "\frac{1}{\lambda} \ln|\cosh(\sqrt{g\lambda }\ t)|"